OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 2y = 4 x – 5y = 11

OpenStudy (johnweldon1993):

Well when you solve that first equation for 'x' you get x = 4 - 2y So after you substitute that into the second equation for 'x' you get 4 - 2y - 5y = 11

OpenStudy (anonymous):

x – 5(2y + 4) = 11 x – 5(–2y + 4) = 11 (2y + 4) – 5y = 11 (–2y + 4) – 5y = 11

OpenStudy (anonymous):

those are my multiple choice options A,B,C,D

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@primeralph helpppp :)

OpenStudy (johnweldon1993):

Oh...just a rearranging of what I put... x = -2y + 4 is the same thing..so -2y + 4 - 5y = 11

OpenStudy (johnweldon1993):

Wait @primeralph is it B? Doesnt the question tell us to isolate 'x' in the first equation?

OpenStudy (johnweldon1993):

No no no...it's D like I said lol

OpenStudy (primeralph):

Oh yeah.

OpenStudy (primeralph):

Need to go eat something.

OpenStudy (anonymous):

Okay thankyou guys :)

OpenStudy (johnweldon1993):

lol go refuel Primeralph! haha...and anytime @juliaann :)